# Label Words are Anchors: An Information Flow Perspective for Understanding In-Context Learning

### v1

@@ Appendix A Experimental Settings
For datasets, we use a sentiment analysis task, Stanford Sentiment Treebank Binary (SST-2) Socher et al. (2013), a question type classification task, Text REtrieval Conference Question Classification (TREC) Li and Roth (2002); Hovy et al. (2001), a topic classification task, AG’s news topic classification dataset (AGNews) Zhang et al. (2015), a emotion classification task, EmoContext (EmoC) Chatterjee et al. (2019). The ICL templates of these tasks are shown in Table 5.
Table 5: Demonstration templates and label words, here <S1> represents the demonstration input, <S> represents the input to be predicted, and <L> represents the label word corresponding to the demonstration output. To save space, we only show one example for each task.
@@
Appendix B Reason for Using Word Loyalty Besides Label Loyalty
Label loyalty alone may not capture changes in the probability of other words or the relative probability distribution of the label words. Word loyalty helps address this limitation, which is shown in Table 6.
Table 6: Results on a test sample with the label ’World’ from AGNews.
@@ Appendix B Reason for Using Word Loyalty Besides Label Loyalty
No isolation World ‘ World’, ‘ Technology’, ‘ Politics’, ‘ Israel’, ‘ Human’
Appendix C Isolating Different Numbers of Layers
We study the impact of the numbers of isolated layers, as shown in Figures 5a and 5b. It can be found that for the first layers, the more layers are isolated, the greater the impact on the model, while the increase in the number of isolated layers has a smaller impact on the model for the latter few layers. This further illustrates the important role of the information aggregation via label words in the shallow layers.
@@ Appendix C Isolating Different Numbers of Layers
Figure 5: Changes in Label Loyalty ($L$) and Word Loyalty (WL) under different numbers of isolated layers. The letter F in brackets indicates the isolation of the label words in the first few layers, and L in brackets indicates the isolation of the label words in the last few layers. The loyalty of isolating the last layers is consistently high, while the loyalty of isolating the first layers decreases as the layer number increases.
Appendix D Results of Isolating Label Words on TREC and EmoC
Relationship between the relative sizes of $S_{wp}$, $S_{pq}$, and $S_{ww}$ and the number of layers on TREC and EmoC are shown in Figure 6.
(a) Results on the TREC dataset
(b) Results on the EmoC dataset
Figure 6: Relationship between the relative sizes of $S_{wp}$, $S_{pq}$, and $S_{ww}$ and the number of layers on AGNews and EmoC.
@@ Appendix E Details for the Calculation of $\text{AUCROC}_{l}$
Suppose the positions of the label words in the input $x$ are $p_{1},...,p_{C}$ (without loss of generality, we suppose $p_{i}$ corresponds to the $i$th class), the targeted position is $q$, the sum of the attention matrices of each attention head at the $l$ layer is $A_{l}$. We postulate that there’s a strong correlation between the attention distributions on the label words of the target position $(A_{l}(q,p_{1}),...,A_{l}(q,p_{C}))$ and the model’s final prediction. We use the AUC-ROC score to quantify this correlation. We regard $(A_{l}(q,p_{1}),...,A_{l}(q,p_{C}))$ as a classifier’s prediction for the model output label (that is, $A_{l}(q,p_{i})$ is equivalent to the probability of model output label $i$), and compute the AUC-ROC value of this prediction relative to the actual model output. We denote this as $\text{AUCROC}_{l}$.
Appendix F Implementation of Anchor Re-Weighting
To implement anchor re-weighting, in the specific computation process of the model, after calculating the attention matrix $A_{l}^{h}$, we multiply each $A_{l}^{h}(q,p_{i})$ by $\exp(\beta^{i}_{lh})$ and continue with the subsequent computations. In other words, for each attention head, we make the following modifications:
@@ Appendix F Implementation of Anchor Re-Weighting
Appendix G The Factor of $L_{\text{demo}}$ and $L_{\mathbf{x}}$
We discuss the factor of the total length of the demonstrations $L_{\text{demo}}$ and the length of the text to be predicted $L_{\mathbf{x}}$ in compression.
@@ Appendix G The Factor of $L_{\text{demo}}$ and $L_{\mathbf{x}}$
GPT2-XL $1.1\times$ $2.5\times$ $1.5\times$ $1.4\times$
GPT-J $1.5\times$ $6\times 2.2$$2\times$ $1.9\times$
$\frac{L_{\text{demo}}+L_{\mathbf{x}}}{L_{\mathbf{x}}}$ $1.9$ $9.7$ $5.1$ $5.4$
@@
From Table 7, we can observe that the larger the ratio of total length to predicted text length, the more significant the efficiency improvement.
Appendix H Calculation of $\hat{\mathbf{k}}$
For the sample sequence $x_{1},...,x_{T}$ to be predicted, we denote the query vectors of the target positions as $\mathbf{q}_{1},...,\mathbf{q}_{T}$. We construct the matrix $\hat{\mathbf{Q}}=(\mathbf{q}_{1}-\overline{\mathbf{q}},...,\mathbf{q}_{T}-\overline{\mathbf{q}})$ by subtracting the mean from the query vectors. Then, we sequentially find the $M$ directions $\mathbf{v}_{1},...,\mathbf{v}_{M}$ with the M largest variation directions for the centralized $\hat{\mathbf{q}_{1}},...,\hat{\mathbf{q}_{T}}$:
@@ Appendix H Calculation of $\hat{\mathbf{k}}$
Additionally, we denote the corresponding standard deviations $\sqrt{\operatorname{Var}\left\{\mathbf{v}_{i}^{\top}\hat{\mathbf{Q}}\right\}}$ as $\sigma_{i}$.
We project the key vector $\mathbf{k}$ onto the directions $\mathbf{v}_{1},...,\mathbf{v}_{M}$ and weight them by $\sigma_{1},...,\sigma_{M}$ to obtain the feature $\hat{\mathbf{k}}$ that reflects the variation of the query vectors. ($\hat{\mathbf{k}}_{i}=\sigma_{i}\mathbf{v}_{i}^{T}\mathbf{k}$)
We study the effect of $M$ in the confusion matrix of $\text{Confusion}_{ij}^{\text{pred}}$ (Figure 7). Since the result under various $M$ is similar, we simply choose $M=10$ to calculate $\text{Confusion}_{ij}^{\text{pred}}$.
@@
Figure 7: Predicted confusion matrixs under $M=5,10,20,50,100,200$.
Appendix I Calculation of $\text{Confusion}_{ij}$
To measure the actual confusion between category $i$ and category $k$ for the model, we propose the metric $\text{Confusion}_{ij}$:
We extract all test samples $x_{t}$ with true labels $i$ or $k$, as well as the probabilities $p_{i}^{t}$ and $p_{j}^{t}$ output by the model for categories $i$ and $k$ on these samples, respectively (normalized to sum up to 1). In other words, we obtain a classifier $f$ that outputs the probabilities $p_{i}^{t}$ and $p_{j}^{t}$ for categories $i$ and $k$, respectively, on the test samples $x_{t}$. We calculate the AUC-ROC value of this classifier $f$ to obtain the degree of confusion between category $i$ and $k$, denoted as $\text{Confusion}_{ij}$. The calculated $\text{Confusion}_{ij}$ is a value that does not exceed 1. The closer $\text{Confusion}_{ij}$ is to 1, the lighter the degree of confusion, and vice versa.
We use the above metric instead of directly analyzing the output labels of the model because previous work has indicated the issue of insufficient output probability calibration in ICL Zhao et al. (2021), which is greatly affected by factors such as sample ordering and model preferences for label words. By using the defined degree of confusion $\text{Confusion}_{ij}$, we can implicitly mitigate the interference of insufficient probability calibration on the output labels, to some extent, obtain the degree of confusion of the model for different categories that better reflects the real situation, and reduce the interference of randomness.

### v2

@@ 5 Conclusion
Limitations
Our study, while providing valuable insights into in-context learning (ICL), has several limitations. Firstly, our research scope was limited to classification tasks and did not delve into the realm of generative tasks. Additionally, our hypothesis was only examined within conventional ICL paradigms, leaving other ICL paradigms such as the chain of thought prompting (CoT) Wei et al. (2022) unexplored. Secondly, due to hardware constraints, we mainly investigated models up to a scale of 6 billion parameters. Further research that replicates our study using larger-scale models would be beneficial in corroborating our findings and refining the hypotheses set forth in our investigation.
Acknowledgement
We thank all reviewers for their thoughtful and insightful suggestions. This work is supported in part by a Tencent Research Grant and National Natural Science Foundation of China (No. 62176002). Xu Sun is the corresponding author.
Appendix
@@ Appendix A Experimental Settings
For datasets, we use a sentiment analysis task, Stanford Sentiment Treebank Binary (SST-2) Socher et al. (2013), a question type classification task, Text REtrieval Conference Question Classification (TREC) Li and Roth (2002); Hovy et al. (2001), a topic classification task, AG’s news topic classification dataset (AGNews) Zhang et al. (2015), and an emotion classification task, EmoContext (EmoC) Chatterjee et al. (2019). The ICL templates of these tasks are shown in Table 4.
Table 4: Demonstration templates and label words. Here <S1> represents the demonstration, <S> represents the input to be predicted, and <L> represents the label word corresponding to the demonstration. To save space, we only show one demonstration for each task.
@@
Appendix B Results of $S_{wp}$, $S_{pq}$, and $S_{ww}$ on TREC and EmoC
Figure 7 illustrates the relative sizes of $S_{wp}$, $S_{pq}$, and $S_{ww}$ on TREC and EmoC, mirroring results on SST-2 and AGNews. In shallow layers, $S_{wp}$ (the information flow from the text part to label words) is prominent, while $S_{pq}$ (the information flow from label words to targeted positions) is less significant. However, in deeper layers, $S_{pq}$ dominates. Importantly, $S_{wp}$ and $S_{pq}$ generally exceed $S_{ww}$, indicating that interactions involving label words are predominant.
(a) Results on the TREC dataset
(b) Results on the EmoC dataset
Figure 7: Relative size of $S_{wp}$, $S_{pq}$, and $S_{ww}$ on TREC and EmoC, which is similar to that on SST-2 and AGNews.
Appendix C Reason for Using Word Loyalty Besides Label Loyalty
@@
No isolation World “ World”, “ Technology”, “ Politics”, “ Israel”, “ Human”
Table 5: Results on a test sample with the label “World” from AGNews.
Label loyalty alone may not capture changes in the probability distribution of non-label words or the relative ratio of the probability of the label words within the entire vocabulary. Word loyalty helps address this limitation, which is shown in Table 5.
Appendix D Isolating Different Numbers of Layers
We study the impact of the numbers of isolated layers, as shown in Figures 8a and 8b. It can be found that isolating shallow layers cause a significant impact, isolating deep layers has a negligible impact on the model, even when the number of isolation layers increases. This further illustrates the important role of information aggregation via label words in the shallow layers.
@@ Appendix D Isolating Different Numbers of Layers
Figure 8: The chart demonstrates variations in label loyalty and word loyalty, dependent on whether label or non-label words are isolated in various layers. ’First’ refers to the first several layers, while ’Last’ to the last ones. Deep-colored lines represent label word isolation, whereas light colors denote non-label words. Remarkably, isolating label words in the shallow layers significantly influences the outcome, regardless of whether this is compared to isolation in deep layers or to non-label word isolation.
@@ Appendix E Details for the Calculation of $\text{AUCROC}_{l}$
Suppose the positions of the label words in the input $x$ are $p_{1},...,p_{C}$ (without loss of generality, we suppose $p_{i}$ corresponds to the $i$th class), the targeted position is $q$, the sum of the attention matrices of all attention heads at the $l$ layer is $A_{l}$. We postulate that there’s a strong correlation between the attention distributions on the label words of the target position $(A_{l}(q,p_{1}),...,A_{l}(q,p_{C}))$ and the model’s final prediction. We use the AUC-ROC score to quantify this correlation. We regard $(A_{l}(q,p_{1}),...,A_{l}(q,p_{C}))$ as a classifier’s prediction for the model output label (that is, $A_{l}(q,p_{i})$ is equivalent to the probability of model outputting label $i$), and compute the AUC-ROC value of this prediction relative to the actual model output. We denote this as $\text{AUCROC}_{l}$. For the case with more demonstrations (Appendix F.1), we simply sum up all $A_{l}(q,p)$ of the same class.
Appendix F Additional Experimental Results
F.1 Results with More Demonstrations
We implement our experimental analysis utilizing two demonstrations per class, resulting in a total of 4, 12, 8, and 8 demonstrations respectively for SST-2, TREC, AGNews, and EmoC. Our findings, as depicted in Figure 9, Figure 10, and Figure 11, exhibit a high degree of similarity to the results obtained from experiments that employ one demonstration per class.
(a) Results on the SST-2 dataset
(b) Results on the TREC dataset
(c) Results on the AGNews dataset
(d) Results on the EmoC dataset
Figure 9: Relative sizes of $S_{wp}$, $S_{pq}$, and $S_{ww}$ when more demonstrations are employed.
(a) Effect of different numbers of isolated layers on GPT2-XL
(b) Effect of different numbers of isolated layers on GPT-J
Figure 10: Variations in label loyalty and word loyalty when more demonstrations are employed.
(a) GPT2-XL (total 48 layers).
(b) GPT-J (total 28 layers).
Figure 11: $\text{AUCROC}_{l}$ and $R_{l}$ of each layer in GPT models when more demonstrations are employed.
F.2 Results for In-Context Learning with semantically-unrelated labels
The applicability of our analytical conclusions to ICL variants, such as the semantically unrelated label ICL Wei et al. (2023), is an intriguing subject. Given that both GPT2-XL and GPT-J-6B perform at levels akin to random guessing in this ICL setting, we chose LLaMA-33B Touvron et al. (2023) and SST-2 for our experiment. We substituted labels with ’A’/’B’, and adhered to a similar experimental setup as in sections § 2.2 and § 2.3. However, we applied eight shots per class to facilitate the model in achieving an accuracy of 83.0% on SST-2. The outcomes align with those derived in § 2.2 and § 2.3. Figure 12 shows the more pronounced impact of isolating labels in the shallow layers compared to their isolation in the deep layers or the isolation of non-label tokens. Figure 13 confirmed that the model leverages information from anchors in the deeper layers to perform classification.
Figure 12: The impact of isolating label words versus randomly isolating non-label words within the first or last 5 layers. Isolating label words within the first 5 layers exerts a more pronounced effect, highlighting the importance of shallow-layer information aggregation via label words.
Figure 13: AUCROCl and $R_{l}$ of each layer of LLaMA-33B on SST-2. Still, deep layers display higher relevance to model prediction, reinforcing the idea that the model extracts information from deep-layer anchors for classification.
Appendix G Implementation of Anchor Re-weighting
In order to implement anchor re-weighting, specific adjustments are made in the model’s computational process. After calculating the attention matrix $A_{l}^{h}$ of the $h$th head in the $l$th layer, we multiply each $A_{l}^{h}(q,p_{i})$ by $\exp(\beta^{i}_{0,lh})$ before proceeding with further computations. This means that for each attention head, we introduce the following modifications:
@@ Appendix G Implementation of Anchor Re-weighting
Appendix H Training Settings of Anchor Re-weighting
For each random seed, we fix the demonstration and sample $1000$ test samples from the test datasets as described in § 2.2. The optimization of parameter vector $\boldsymbol{\beta}$ is carried out using gradient descent, specifically with the Adam optimizer Kingma and Ba (2015). The learning rate is set at $0.01$, with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. Due to memory constraints, we use a batch size of 1. This optimization process is repeated for $10$ epochs. Owing to limitations in computational resources, we restrict our evaluation to the GPT2-XL model and exclude the GPT-J model from our assessment.
Appendix I The Factor of $L_{\text{demo}}$ and $L_{\mathbf{x}}$
@@
GPT2-XL $1.1\times$ $1.5\times$ $2.5\times$ $1.4\times$
GPT-J $1.5\times$ $2.2\times$ $2.9\times$ $1.9\times$
$L_{\text{demo}}$ $18$ $61$ $151$ $53$
$L_{\mathbf{x}}$ 1919 77 $37$ $12$
Table 6: Acceleration ratios, $L_{\text{demo}}$ and $L_{\mathbf{x}}$.
From Table 6, we observe a correlation between the acceleration ratios and the ratio of the total demonstration length ($L_{\text{demo}}$) to the length of the text predicted ($L_{\mathbf{x}}$). It suggests that a greater ratio of total length to predicted text length may yield a higher acceleration ratio.
@@
Appendix J Calculation of $\hat{\mathbf{k}}$
For the sampled sequence $x_{1},...,x_{T}$ to be predicted, we denote the query vectors of the target positions as $\mathbf{q}_{1},...,\mathbf{q}_{T}$. We then compute the matrix $\hat{\mathbf{Q}}=(\mathbf{q}_{1}-\overline{\mathbf{q}},...,\mathbf{q}_{T}-\overline{\mathbf{q}})$ by subtracting the mean vector, $\overline{\mathbf{q}}$, from each query vector. Subsequently, we determine the $M$ directions, $\mathbf{v}_{1},...,\mathbf{v}_{M}$, that correspond to the M largest variation directions for the centralized query vectors $\hat{\mathbf{q}}_{1},...,\hat{\mathbf{q}}_{T}$. The $i^{th}$ direction, $\mathbf{v}_{i}$, is chosen to maximize the variance of the projection of the centralized query vectors onto it, while also being orthogonal to the previously chosen directions, $\mathbf{v}_{1},...,\mathbf{v}_{i-1}$. This process can be formalized as follows:
@@
We define $\sigma_{i}$ as the square root of the variance of the projection of $\hat{\mathbf{Q}}$ onto the $i^{th}$ direction, i.e., $\sqrt{\operatorname{Var}\left\{\mathbf{v}_{i}^{\top}\hat{\mathbf{Q}}\right\}}$.
To derive features $\hat{\mathbf{k}}$s, we project the key vector $\mathbf{k}$ onto the directions $\mathbf{v}_{1},...,\mathbf{v}_{M}$ and scale the projections by the corresponding standard deviations $\sigma_{1},...,\sigma_{M}$. Each feature, $\hat{\mathbf{k}}_{i}$, is thus calculated as $\sigma_{i}\mathbf{v}_{i}^{T}\mathbf{k}$.
We further examine the influence of $M$ on the prediction confusion matrix, $\text{Confusion}{ij}^{\text{pred}}$, as depicted in Figure 14. Given the similarity in outcomes for various $M$, we settle on a value of $M=10$ for computation of $\text{Confusion}{ij}^{\text{pred}}$.
@@
Figure 14: Predicted confusion matrices under $M=5,10,20,50,100,200$.
Appendix K Calculation of $\text{Confusion}_{ij}$
To gauge the true degree of confusion between categories $i$ and $k$ for a given model, we suggest utilizing the $\text{Confusion}_{ij}$ metric:
First, we procure all test samples $x_{t}$ bearing true labels $i$ or $k$. We then obtain the probabilities $p_{i}^{t}$ and $p_{j}^{t}$ yielded by the model for categories $i$ and $k$, respectively, on these samples. These probabilities are normalized to a total of 1. Essentially, we derive a classifier $f$ that delivers the probabilities $p_{i}^{t}$ and $p_{j}^{t}$ for the categories $i$ and $k$ respectively, on the test samples $x_{t}$. By calculating the Area Under the Receiver Operating Characteristic Curve (AUC-ROC) value of this classifier $f$, we get the degree of confusion between category $i$ and $k$, termed as $\text{Confusion}_{ij}$.
The computed $\text{Confusion}{ij}$ is a value that never exceeds 1. The closer $\text{Confusion}{ij}$ approximates 1, the less pronounced the confusion, and vice versa.
We use the above metric instead of directly analyzing the output labels of the model because previous work has indicated the issue of insufficient output probability calibration in ICL Zhao et al. (2021), which is greatly affected by factors such as sample ordering and model preferences for specific label words. By leveraging our defined degree of confusion, $\text{Confusion}_{ij}$, we can implicitly alleviate the disturbances arising from insufficient probability calibration on the output labels. This allows for a more accurate representation of the model’s degree of confusion for different categories, mitigating the impact of randomness.
Appendix L Reproducibility
In the supplementary material, we have provided codes that allow for the faithful replication of our experiments and subsequent result analysis. To ensure consistency and reproducibility across different devices, we have fixed the five random seeds to the values of 42, 43, 44, 45, and 46. We invite readers to delve into the code for additional implementation details that may arouse their interest.